Write the formula for the complex ion formed by the metal ion Cr3+ and six NO2- ions as ligands. Decide whether the complex could be isolated as a chloride salt or a potassium salt, and write the formula for the appropriate salt
Answer #1

Cr has +3 charge. NO2 has -1 charge

we have 6 NO2. so, total negative charge = 6*(-1) = -6

On combining total charge = +3 -6 = -3

Thus, complex is [Cr(NO2)6]3-

Since the complex has negative charge. So it will be combined with positive ion i.e. K+ to balance the charge.Thus, the complex could be isolated as a potassium salt.

and the formula for the salt is K3 [Cr(NO2)6]